# AI Agentic Programming: A Survey of Techniques, Challenges, and Opportunities

### v1

@@ 6. Opportunities and Future Directions
AI agentic programming represents a fast-evolving research frontier that intersects artificial intelligence, programming languages, and software engineering. While recent advances have demonstrated promising capabilities, significant challenges remain in realizing robust, efficient, and trustworthy agentic systems. In this section, we outline several key opportunities and open research directions that can shape the future of this field.
@@ 6.1. Integrating Coding Agents with Tools
A promising direction is to rethink how programming languages, compilers, and testing frameworks, which are traditionally built for human developers, can be redesigned to support AI coding agents. For example, instead of emitting opaque diagnostics, compilers could provide structured feedback explaining why certain optimizations (e.g., vectorization or inlining) fail (maleki2011evaluation, ; ashouri2018survey, ; siso2019evaluating, ). These could include semantic barriers like unresolved aliasing, ambiguous data/control flow, or missing annotations (bi2024iterative, ), enabling agents to revise code more precisely. Beyond diagnostics, compilers can help agents track state across iterations. Feedback on which edits introduced errors, failed assertions, or performance regressions would enable agents to reason over the change history and adjust strategies accordingly.
Opening compiler internals also presents a valuable opportunity. Coding agents could interact directly with intermediate representations (IRs), such as LLVM IR (lattner2004llvm, ) or MLIR (lattner2021mlir, ), to reason about program structure, verify transformations, or perform static analysis at a semantic level. Compiler APIs and language servers (e.g., Clang’s LibTooling, the Language Server Protocol) already expose ASTs, symbol tables, and refactoring tools, but a wider adoption may require standardized, introspective interfaces across compilers.
@@
Tighter integration with runtime systems also offers opportunities. For instance, agents can dynamically insert instrumentation or launch profiling runs, then use the results to inform optimization choices. Coupling these capabilities with autotuning frameworks (chen2018tvm, ; ansel2014opentuner, ) would expand the design space while preserving correctness and safety.
Finally, advances in structured code representations, such as ASTs, graph-based IRs, and semantic embeddings, offer a foundation for more powerful agent reasoning. Combining LLMs with graph neural networks or neuro-symbolic systems could improve generalization and support cross-language, cross-target understanding.
@@ 6.2. Scalable Memory and Context Management
A key capability of agentic programming lies in managing memory and contextual information across tasks that involve long context reasoning and multiple iterations. Unlike traditional code generation, which typically follows a single pass prompt to solution model, agentic workflows for solving real-world software engineering problems involve multiple steps, iterative refinement, and integration with external tools and development environments (cursor, ; wang2024openhands, ; openai2021codex, ; github_copilot, ).
@@
As the memory footprint of LLMs grows linearly with input token length (DBLP:conf/nips/VaswaniSPUJGKP17, ; kaplan2020scaling, ), current LLM-based agents remain constrained by their context windows and lack persistent memory across a long sequence of iterations (schick2023toolformer, ; liu2023lost, ; packer2023memgpt, ). However, reasoning about real-world programs often requires modeling complex data structures and code context (like function calls) spanning across multiple files, which often exceeds typical context limits. Although some industry-scale LLMs claim to support million-token contexts (ding2023longnet, ; team2024gemini, ; ding2024longrope, ), they often rely on random sampling techniques (hosseini2024efficient, ) and fail to leverage program structure or semantics effectively.
Therefore, an interesting direction is to design attention mechanisms that are guided by code structure, such as syntax trees, control flow graphs, or data dependencies. These structures can help agents focus more accurately on the most relevant parts of a program. While approaches like retrieval-augmented generation (RAG) (lewis2020retrieval, ), KV cache offloading (lee2024infinigen, ; sun2024shadowkv, ), and compression (liu2024kivi, ; chang2025palu, ) provide partial solutions, they struggle to provide precise control over long-term dependencies, structured knowledge, and execution histories.
@@
Another important area is the development of context-aware retrieval strategies that move beyond static similarity-based methods (su2024dragin, ). During debugging, for instance, an agent could retrieve not only the most recent error message but also similar past failures, proposed fixes, relevant test cases, and their outcomes. Retrieval conditioned on task state and tool feedback would significantly improve the agent’s ability to reason under uncertainty.
@@ 6.3. Evaluation and Benchmarking
Abbreviations: CP = Competitive Programming, BF = Bug Fixing, FC = Function Completion, CR = CLI Reasoning.
Benchmark Source Language Task Difficulty Year
HumanEval Hand-written Python FC Beginner 2021
MBPP Crowd-sourced Python FC Beginner 2021
CodeContests CP Python/C++/Java FC Diverse 2022
HumanEval-X Hand-written Python/C++/Java/JS/Go FC Intermediate 2023
SWE-Bench GitHub Issues Python BF Expert 2024
SWE-bench M GitHub Issues JS BF Diverse 2024
LiveCodeBench CP (live) Python FC, BF Diverse 2024
TerminalBench Community-curated Shell CR Diverse 2025
Spider 2.0 Enterprise DB Apps SQL FC Expert 2025
EffiBench-X Synthetic Python/C++/JS FC, BF Diverse 2025
Web-Bench Web App Projects JS/TS/HTML/CSS/Python BF, FC Expert 2025
ProjectEval Open-source repos Python/Java/C++/JS FC, BF, CR Expert 2025
TRAIL CP Python/Java/C++/JS BF, CR Expert 2025
@@
Table 7. AI Intelligence Score (across mathematics, science, coding, and reasoning) and Speed (Output tokens per second) for LLMs on LivecodeBench
Metric GPT-5 GPT-4o Claude 4 Gemini 2.5 Pro Grok 4 DeepSeek R1 Kimi K2 Qwen3 Solar-Pro Devstral Llama 4 Magistral
AI score 69 65 59 65 68 59 49 64 43 31 42 36
Speed 102 148 100 153 76 21 45.2 24 54 44 174 51
@@
Table 6 summarizes several widely used coding benchmarks for evaluating LLMs and agentic systems on programming tasks. These benchmarks vary in their origin, programming language coverage, task types, and difficulty levels. Commonly used datasets include HumanEval (openai2021codex, ), HumanEval-X (zheng2023codegeex, ), MBPP (austin2021program, ), SWE-Bench (jimenez2023swe, ), SWE-bench Multimodal (yang2024swe, ), TerminalBench (tbench_2025, ), LiveCodeBench (jain2024livecodebench, ), CodeContests (DBLP:journals/corr/abs-2203-07814, ), Spider2.0 (lei2024spider, ), EffiBench-X (qing2025, ), Web-Bench (xu2025webbench, ), ProjectEval (liu-etal-2025-projecteval, ), and TRAIL (deshpande2025trailtracereasoningagentic, ).
@@
While these benchmarks have provided valuable insights into the capabilities of LLMs and agentic systems for code generation and bug fixing, they also have important limitations. For example, they are heavily biased toward a small set of programming languages - especially Python, which dominates the training corpus of code for current models (twist2025llms, ; chai2024mceval, ). This limits the generalizability of evaluation results to domains involving statically typed or domain-specific languages, such as C++ or Rust (twist2025llms, ). For example, Table 7 presents the AI Intelligence Score (aggregated across mathematics, science, coding, and reasoning) and the output speed (in tokens per second) for a selection of widely used LLMs on the LiveCodeBench evaluation. For the intelligence score, the latest model, GPT-5, achieved 69, indicating that more than 30% of challenges remain unresolved.
@@
Furthermore, many existing benchmarks focus on small, self-contained problems that may not be representative of real-world software engineering tasks. Realistic development scenarios often involve working with large, modular codebases, extensive use of third-party libraries, non-trivial build processes, and long-range dependencies across files and components. These characteristics are largely absent from current benchmark datasets, making it difficult to assess an agent’s ability to scale or generalize.
Another key limitation is that most benchmarks do not capture the interactive, iterative nature of agentic programming. In real-world settings, coding agents will need to collaborate with human developers (wang2024openhands, ; cursor, ; openai2021codex, ; github_copilot, ), receive intermediate feedback and confirmation, and rely on external tools such as compilers, debuggers, and test frameworks. Benchmarks that assume single-shot or non-interactive task completion fail to reflect the complexity of such multi-step, tool-augmented workflows.
@@
Addressing these gaps will require the development of more comprehensive and extensible evaluation frameworks. Future benchmarks should incorporate realistic tasks that reflect end-to-end development workflows, support multiple programming languages, and enable interaction with tools and human feedback loops. Metrics should go beyond functional correctness to include robustness, tool usage efficiency, recovery from failure, and the ability to incorporate feedback. Simulation environments and evaluation harnesses will also be important for reproducibility and fair comparison.
@@ 6.4. Human-AI Collaboration
While a long-term vision of agentic programming is to automate the entire software development lifecycle, including writing, debugging, and testing code, near-term opportunities include extending the capabilities of current LLM-based coding assistants (wang2024openhands, ; cursor, ; openai2021codex, ; github_copilot, ). Rather than replacing human developers, these systems can act as collaborative partners, supporting workflows in which humans retain strategic oversight. Similar to pair programming (hannay2009effectiveness, ), LLM-based agents can assist by proposing ideas, detecting errors, suggesting improvements, and automating routine tasks, thereby augmenting human productivity.
@@
Designing effective human–AI collaboration models remains an important research challenge. This includes developing user interfaces, interaction protocols, and frameworks for allocating responsibilities between human developers and agents, particularly in professional and team-based environments where coordination, trust, and efficiency are essential. Unlike traditional code generation tools, agentic systems operate in iterative, feedback-driven loops and are capable of making autonomous decisions based on intermediate results. This interaction model shifts the role of the developer from directive command-giver to interactive collaborator, opening up new opportunities and challenges for co-creating software. To enable meaningful collaboration, agents must be transparent in their reasoning, responsive to human input, and able to explain their decisions. Realizing these capabilities will require advances in interactive prompting (schulhoff2024prompt, ), natural language explanations (creswell2022faithful, ), and context-aware dialogue protocols (shah2025prompt, ; anthropic2024mcp, ).
One promising research direction is the development of mixed-initiative workflows, in which control shifts fluidly between the developer and the agent. In these workflows, the human might specify high-level goals, architectural constraints, or coding conventions, while the agent proactively generates scaffolding code, explores design alternatives, or automates repetitive subtasks. The developer can then inspect, accept, reject, or refine the agent’s contributions. For instance, an agent could propose multiple refactoring strategies for a large function and generate test cases to verify the correctness of each variant. The human collaborator evaluates these options and selects the most appropriate one, potentially with further refinements. Supporting such workflows requires robust management of shared context, memory, and goal state across multiple interaction turns.
Agents must also be able to handle ambiguity and uncertainty. Real-world development often involves vague or underspecified requirements. Future research may explore agents that can ask clarifying questions (zhang2024modeling, ; mu2024clarifygpt, ), infer intent from surrounding artifacts, or learn user preferences over time. For example, when given an instruction like “make this more efficient”, the agent should determine whether to optimize for speed, memory, or readability - potentially by prompting the user for clarification.
Furthermore, human developers rely on a rich ecosystem of non-code artifacts, such as design diagrams, issue trackers, documentation, and domain-specific knowledge. Integrating these contextual signals into the agent’s reasoning pipeline remains an open challenge. For instance, when fixing a bug, the agent may need to cross-reference issue tracker discussions or past commits to fully understand the context and constraints of the fix.
Trust and usability are foundational to collaboration. Developers must feel confident that the agent’s actions are safe, interpretable, and reversible. Research is needed on techniques that allow agents to explain their decisions, cite relevant documentation or examples (e.g., from internal codebases or public repositories), and highlight trade-offs. Users should be empowered to inspect, override, or revert agent-generated changes easily—especially in safety-critical or compliance-sensitive domains such as finance, healthcare, or aerospace.
Collaboration in team-based environments presents additional opportunities. Future systems may support multi-user workflows where agents assist in coordinating tasks across a team, help onboard new developers by answering project-specific questions, or even mediate merge conflicts by suggesting consistent resolutions based on prior history. For example, an agent integrated into a version control system might suggest commit messages, flag inconsistent code style across pull requests, or recommend reviewers based on code ownership.
In summary, enabling rich human-AI collaboration in agentic programming will require progress across multiple dimensions: user interface design, communication strategies, intent modeling, memory management, and trust calibration. These systems should be designed not to replace human developers, but to augment their creativity, efficiency, and decision-making. Ultimately, effective collaboration will depend on how well agents can adapt to the social and cognitive workflows of modern software development.
@@ 6.5. Domain Specialization and Adaptability
General-purpose coding agents often underperform in domain-specific tasks due to a lack of tailored knowledge, tool integration, and contextual understanding (joel2024survey, ; zheng2024well, ). While LLMs trained on extensive code corpora can demonstrate general programming capabilities, they may struggle in specialized domains, such as generating Verilog code for hardware synthesis, which involves strict performance constraints, low-level abstractions, or domain-specific libraries and tooling.
Future research may explore strategies for domain-adaptive prompting, fine-tuning, and agent behavior modulation to improve performance in areas such as embedded systems, data science, scientific computing, high-performance computing, and formal methods. For example, in embedded systems programming (englhardt2024exploring, ), agents must reason about memory layout, real-time constraints, and hardware-specific APIs, which are often underrepresented in general training datasets of code. Fine-tuning on domain-specific codebases or adapting the agent’s planning strategy to prioritize correctness and safety over code brevity may significantly improve performance. Another direction is to build LLM agents that incorporate domain-specific tools and diagnostics into their reasoning loop. For instance, a scientific computing agent might integrate with tools like NumPy, MATLAB, or domain-specific simulation engines (openfoam, ; permann2020moose, ; simulink, ), allowing it to generate, run, and analyze numerical experiments autonomously. Similarly, an agent for cybersecurity tasks might interact with static analyzers (codeql2020, ), fuzzers (manes2019art, ), or formal verification tools (woodcock2009formal, ) to check code properties as part of its development process.
Domain specialization can also improve interpretability and safety. By narrowing the operational scope of an agent, developers can apply stricter validation, leverage existing formal specifications, and generate more meaningful feedback. In safety-critical applications, such as aerospace, automotive, or medical software, agents must operate under rigorous constraints, and tailoring their reasoning and code generation behaviors to domain-specific requirements is essential.

### v2

@@ 6. Opportunities and Future Directions
Software Compatibility (Section 6.1) Structured feedback on failures Direct IR interaction Standardized compiler interfaces How do coding agents integrate with tools? Context Management (Section 6.2) Structured storage and recall of evolving task information Conditioned retrieval from task state and tool feedback Program state tracing and replay How do coding agents manage memory and contextual information? Benchmarks (Section 6.3) Heavily biased training corpus Large absence of real-world tasks Lack of reflecting interactive and iterative programming workflow How to comprehensively evaluate coding agents? Human-AI Collaboration (Section 6.4) Frameworks for responsibility allocation Handle uncertainty; Ecosystem integration Coordination in multi-user environments How do coding agents collaborate with human? Domain-specific Agents (Section 6.5) Adaptive prompting, fine-tuning, and behavior modulation Reasoning-loop integration of domain tools and diagnostics Runtime domain detection and adaptation How to develop domain-specific coding agents? Safety, Alignment and Trust (Section 6.6) Agent behavior grounding Risk detection and reduction Explainable and self-reflective agents How to ensure safer, ethical, and trustworthy coding agents? Multi-Agent Collaboration (Section 6.7) Native protocol support for multi-agent systems Session management for dynamic coordination Semantics-aware routing, conflict resolution, and consistency How to design standardized protocols for multi-agent collaboration? System Support (Section 6.8) System resource scheduling (CPU, GPU, memory) Reducing overhead in communication with environments Coordinating tasks and resolving conflicts in collaborative codebases How to enhance system-level support for coding agents?
Figure 8. Summary of potential future research directions of AI agentic programming.
AI agentic programming represents a fast-evolving research frontier that intersects artificial intelligence, programming languages, and software engineering. While recent advances have demonstrated promising capabilities, significant challenges remain in realizing robust, efficient, and trustworthy agentic systems. In this section, we outline several key opportunities and open research directions that can shape the future of this field. An overview is provided in Figure 8.
@@ 6.1. Integrating Coding Agents with Tools
A promising direction is to rethink how programming languages, compilers, and testing frameworks, which are traditionally built for human developers, can be redesigned to support AI coding agents. For example, instead of emitting opaque diagnostics, compilers could provide structured feedback explaining why certain optimizations (e.g., vectorization or inlining) fail (Maleki et al., 2011a; Ashouri et al., 2018; Siso et al., 2019). These could include semantic barriers like unresolved aliasing, ambiguous data/control flow, or missing annotations (Bi et al., 2024), enabling agents to revise code more precisely. Beyond diagnostics, compilers can help agents track state across iterations. Feedback on which edits introduced errors, failed assertions, or performance regressions would enable agents to reason over the change history and adjust strategies accordingly.
Opening compiler internals also presents a valuable opportunity. Coding agents could interact directly with intermediate representations (IRs), such as LLVM IR (Lattner and Adve, 2004b) or MLIR (Lattner et al., 2021), to reason about program structure, verify transformations, or perform static analysis at a semantic level. Compiler APIs and language servers (e.g., Clang’s LibTooling, the Language Server Protocol) already expose ASTs, symbol tables, and refactoring tools, but a wider adoption may require standardized, introspective interfaces across compilers.
@@
Tighter integration with runtime systems also offers opportunities. For instance, agents can dynamically insert instrumentation or launch profiling runs, then use the results to inform optimization choices. Coupling these capabilities with autotuning frameworks (Chen et al., 2018a; Ansel et al., 2014) would expand the design space while preserving correctness and safety. Finally, advances in structured code representations, such as ASTs, graph-based IRs, and semantic embeddings, offer a foundation for more powerful agent reasoning. Combining LLMs with graph neural networks or neuro-symbolic systems could improve generalization and support cross-language, cross-target understanding.
@@ 6.2. Scalable Memory and Context Management
A key capability of agentic programming lies in managing memory and contextual information across tasks that involve long context reasoning and multiple iterations. Unlike traditional code generation, which typically follows a single pass prompt to solution model, agentic workflows for solving real-world software engineering problems involve multiple steps, iterative refinement, and integration with external tools and development environments (cur, 2025; Wang et al., 2024d; OpenAI, 2021; GitHub, 2021).
@@
As the memory footprint of LLMs grows linearly with input token length (Vaswani et al., 2017; Kaplan et al., 2020), current LLM-based agents remain constrained by their context windows and lack persistent memory across a long sequence of iterations (Schick et al., 2023; Liu et al., 2023b; Packer et al., 2023). However, reasoning about real-world programs often requires modeling complex data structures and code context (like function calls) spanning across multiple files, which often exceeds typical context limits. Although some industry-scale LLMs claim to support million-token contexts (Ding et al., 2023; Team et al., 2024; Ding et al., 2024), they often rely on random sampling techniques (Hosseini et al., 2024) and fail to leverage program structure or semantics effectively.
Therefore, an interesting direction is to design attention mechanisms that are guided by code structure, such as syntax trees, control flow graphs, or data dependencies. These structures can help agents focus more accurately on the most relevant parts of a program. While approaches like retrieval-augmented generation (RAG) (Lewis et al., 2020), KV cache offloading (Lee et al., 2024; Sun et al., 2024a), and compression (Liu et al., 2024g; Chang et al., 2025) provide partial solutions, they struggle to provide precise control over long-term dependencies, structured knowledge, and execution histories.
@@
Another important area is the development of context-aware retrieval strategies that move beyond static similarity-based methods (Su et al., 2024). During debugging, for instance, an agent could retrieve not only the most recent error message but also similar past failures, proposed fixes, relevant test cases, and their outcomes. Retrieval conditioned on task state and tool feedback would significantly improve the agent’s ability to reason under uncertainty.
@@ 6.3. Evaluation and Benchmarking
Abbreviations: CP = Competitive Programming, BF = Bug Fixing, FC = Function Completion, CR = CLI Reasoning, PO = Performance Optimization.
[…]
@@
Table 8 summarizes several widely used coding benchmarks for evaluating LLMs and agentic systems on programming tasks. These benchmarks vary in their origin, programming language coverage, task types, and difficulty levels. Commonly used datasets include HumanEval (OpenAI, 2021), HumanEval-X (Zheng et al., 2023b), MBPP (Austin et al., 2021b), SWE-Bench (Jimenez et al., 2023), SWE-bench Multimodal (Yang et al., 2024b), Terminal-Bench (Team, 2025), LiveCodeBench (Jain et al., 2024), CodeContests (Li et al., 2022), Spider2.0 (Lei et al., 2024), EffiBench-X (Qing et al., 2025), Web-Bench (Xu et al., 2025b), ProjectEval (Liu et al., 2025), TRAIL (Deshpande et al., 2025), and GSO (Shetty et al., 2025).
@@
While these benchmarks have provided valuable insights into the capabilities of LLMs and agentic systems for code generation and bug fixing, they also have important limitations. For example, they are heavily biased toward a small set of programming languages - especially Python, which dominates the training corpus of code for current models (Twist et al., 2025; Chai et al., 2024). This limits the generalizability of evaluation results to domains involving statically typed or domain-specific languages, such as C++ or Rust (Twist et al., 2025). For example, Figure LABEL:fig:res presents the performance of the top-10 LLMs across four representative benchmarks. The results reveal clear task-dependent differences. For example, all models still perform poorly on software optimization tasks, while they perform well on code generation tasks. Overall, performance gaps remain across all benchmarks, indicating that current LLMs still face significant challenges in reasoning, coding, and generalization to complex tasks.
@@
Furthermore, many existing benchmarks focus on small, self-contained problems that may not be representative of real-world software engineering tasks. Realistic development scenarios often involve working with large, modular codebases, extensive use of third-party libraries, non-trivial build processes, and long-range dependencies across files and components. These characteristics are largely absent from current benchmark datasets, making it difficult to assess an agent’s ability to scale or generalize. Another key limitation is that most benchmarks do not capture the interactive, iterative nature of agentic programming. In real-world settings, coding agents will need to collaborate with human developers (Wang et al., 2024d; cur, 2025; OpenAI, 2021; GitHub, 2021), receive intermediate feedback and confirmation, and rely on external tools such as compilers, debuggers, and test frameworks. Benchmarks that assume single-shot or non-interactive task completion fail to reflect the complexity of such multi-step, tool-augmented workflows.
@@
Addressing these gaps will require developing more comprehensive and extensible evaluation frameworks. Future benchmarks should incorporate realistic tasks that reflect end-to-end development workflows, support multiple programming languages, and enable interaction with tools and human feedback loops. Metrics should go beyond functional correctness to include robustness, tool usage efficiency, recovery from failure, and the ability to incorporate feedback. Simulation environments and evaluation harnesses will also be important for reproducibility and fair comparison.
@@ 6.4. Human-AI Collaboration
While a long-term vision of agentic programming is to automate the entire software development lifecycle, including writing, debugging, and testing code, near-term opportunities include extending the capabilities of current LLM-based coding assistants (Wang et al., 2024d; cur, 2025; OpenAI, 2021; GitHub, 2021). Rather than replacing human developers, these systems can act as collaborative partners, supporting workflows in which humans retain strategic oversight. Similar to pair programming (Hannay et al., 2009), LLM-based agents can assist by proposing ideas, detecting errors, suggesting improvements, and automating routine tasks, thereby augmenting human productivity.
@@
Designing effective models for human-AI collaboration is a key research challenge. This involves creating user interfaces, interaction protocols, and responsibility-sharing frameworks suited for professional and team-based environments where coordination, trust, and efficiency are critical. Unlike traditional code generation tools, agentic systems operate in iterative, feedback-driven loops and make autonomous decisions based on intermediate results. This shifts the developer’s role from command giver to collaborator, opening new opportunities and challenges for co-creating software. To support this shift, agents must be transparent in their reasoning, responsive to human input, and capable of explaining their decisions, which calls for advances in interactive prompting (Schulhoff et al., 2024), natural language explanations (Creswell and Shanahan, 2022), and context-aware dialogue protocols (Shah, 2025; Anthropic, 2024a).
An interesting direction would be mixed initiative workflows, where control moves fluidly between human and agent. Developers can specify goals, constraints, or coding conventions, while agents generate scaffolding code, explore alternatives, or automate repetitive tasks. The developer then inspects, accepts, rejects, or refines the output. For example, an agent may suggest several refactoring options and generate test cases to validate them, leaving the final choice to the human. Such workflows must also support interruption and resumption, enabling agents to incorporate new input without losing context. This requires toolchains that track shared state across iterations, including goals, partial programs, and transformation history.
Finally, domain-aware collaboration can strengthen interpretability and safety. Agents should handle ambiguous or incomplete requirements by asking clarifying questions (Zhang et al., 2024c; Mu et al., 2024), drawing on design documents, issue trackers, and domain knowledge to refine their reasoning. They should also provide clear explanations, cite relevant evidence, and highlight trade-offs to build trust, which is especially critical in domains such as finance, healthcare, and aerospace.
@@ 6.5. Domain Specialization and Adaptability
While LLMs trained on extensive code corpora can demonstrate general programming capabilities (Joel et al., 2024; Zheng et al., 2024a), they may struggle in specialized domains, such as generating Verilog code for hardware synthesis, which involves strict performance constraints, low-level abstractions, or domain-specific libraries and tooling.
Future research can improve agentic programming by exploring domain adaptive prompting, fine-tuning, and behavior modulation in areas such as embedded systems, data science, scientific computing, high-performance computing, and formal methods. In embedded systems programming (Englhardt et al., 2024), for example, agents must reason about memory layout, timing constraints, and hardware-specific APIs, which are often underrepresented in general training data. Training on domain-specific code or adjusting planning strategies to prioritise correctness and safety over brevity may enhance performance. Integrating domain-specific tools and diagnostics into agents’ reasoning, such as simulation engines (OpenFOAM Foundation, 2023; Permann et al., 2020; The MathWorks, Inc., 2023), static analyzers (Lab, 2020), fuzzers (Manès et al., 2019), or formal verification tools (Woodcock et al., 2009), can further improve robustness. Specialisation also supports interpretability and safety, enabling stricter validation, better use of formal specifications, and more reliable feedback in safety-critical domains such as aerospace, automotive, and healthcare.